A.) You have an item you don't want to run out of, but it is not a critical part of your sales. You would like to serve 98% of your customers when they want this item. Looking back through your sales history you find the standard deviation of the product is 11. You also find the EOQ is 20. What is your required safety stock to maintain this level of customer service?

B.) Your sales become more consistent with this product and you now have a standard deviation of 7. Now how much of your inventory is safety stock?
Answer #1

Safety stock = zvalue*standard deviation

For 98% customer service, z value using NORMSINV() function in excel or z table is 2.05

A) For standard deviation = 11

Safety stock = 2.05*11 = 22.55

B) For standard deviation = 7

Safety stock = 2.05*7 = 14.35